Pine Company Ltd. is planning to market a new product, Pinew. To finance the venture, it proposes to have a rights issue at Shs 40/= of one new share for each two shares held. The company currently has 2,000,000 shares outstanding with a market price of Shs 150/= per share. Assume that the new funds are invested to earn the same rate of return as Pine’s other assets.

Required:

i.              Share price after 100% subscription of the issue.                               (6 marks)

ii.            Value of the right to buy one share.                                                   (3 marks)
Answer #1

As we know that the formula for Theroetical Ex Rights price is given by

New Shares * Issue Price + Old Shares * Market Price / New Shares + Old Shares

Now we know that the rights is given for one share for every two shares thus the 100% subscription is done when 1,000,000 shares are subscribed as the old shares are total 2,000,000 (1 for every 2 so for 1 million for all 2 million shares)

Now the Market Price = 150

New Shares = 1,000,000

Subscription Price = 40

Old Shares = 2,000,000

= (1000000*40)+(2000000*150)/(1000000+2000000)

= 40,000,000 + 300,000,000/3,000,000

= 340,000,000/3,000,000

= 113.33Sh

Value of Right = (Stock Price - Rights Price)/(No of Shares required to buy rights + no of right shares)

= (150-40)/3

= 36.667Sh